# Distribution estimation via Flow Matching with Lipschitz guarantees

### v1

@@ 3 Wasserstein distance in Flow Matching
(Kunkel and Trabs, 2025b, Theorem 3.1) Assume that all functions in $\mathcal{M}$ are Lipschitz continuous for fixed $t$ with Lipschitz constant $\Gamma_{t}.$ Then for any $\tilde{v}\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d},$ with $\tilde{v}\in\mathcal{M}$ and $v$ from (6)
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{\psi}_{1}(Z)})\leq\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi_{1}(Z)})+\sqrt{2e}e^{\int_{0}^{1}\Gamma_{t}\;\mathrm{d}t}\Big{(}\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ X_{t}\sim p_{t}\end{subarray}}\big{[}|v_{t}(X_{t})-\tilde{v}_{t}(X_{t})|^{2}\big{]}\Big{)}^{1/2}.$ (11)
The first term in (11) is determined solely by the conditional vector fields $v_{t}(\cdot|y)$and is therefore independent of the approximation class $\mathcal{M}$. To bound this term, we impose the following assumptions on the densities of the unknown distribution $\mathbb{P}^{*}$ and the latent distribution $\mathbb{U}$.
@@
    $\int yp_{0}(y)\;\mathrm{d}y=0,\quad\text{and}\quad\int|y|^{1+\alpha}p_{0}(y)\;\mathrm{d}y<\infty.$
Under ˜3.2, we know that
see proof of (Kunkel and Trabs, 2025b, Theorem 3.3), where the first part of the proof does not use the assumption that the support of $\mathbb{P}^{*}$ is compact. Note that the constant in (12) depends only on $\|p^{*}\|_{B^{\alpha}_{1,\infty}}$ and $d$.
The second term in (11) is connected to the suitability of the set $\mathcal{M}$ to approximate $v$. This depends on the specific structure of $v$. The factor $\exp(\int_{0}^{1}2\Gamma_{t}\;\mathrm{d}t)$ results from the use of Grönwall’s inequality and is one of the key difficulties in the analysis of Flow Matching. In view of Equation˜4, any Lipschitz constraint on the class $\mathcal{M}$ should be aligned with the Lipschitz constant of $v.$ Looking at (6), this Lipschitz constant depends on the construction of $v_{t}(\cdot|\cdot)$ and $p_{t}(\cdot|\cdot)$ and the unknown density $p^{*}$.
In the next section, we are going to study this Lipschitz constant. We are going to focus on one specific probability path and assume the following:
Assumption 3.3.
@@
where the mean shift $\mu\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ and the variance function $\sigma\colon[0,1]\rightarrow\mathbb{R}_{>0}$ are smooth, monotone functions such that
$\mu_{0}(y)=0,$ $\quad\mu_{1}(y)=y,$
$\sigma_{0}=1,$ $\quad\sigma_{1}=\sigma_{\min},$
for a $\sigma_{\min}>0$.
˜3.3 means in particular that the latent distribution is $\mathcal{N}(0,I_{d})$ The monotonicity assumption guarantees that there is no unnecessary movement of mass. ˜3.3 coincides with the construction in Lipman et al. (2023, Section 4). Very similar generative models based on probability Flows have been studied by Gao et al. (2024b, a). Additionally, the use of $\mathcal{N}(0,I_{d})$ for the latent distribution is pervasive in generative modeling, e.g. in Diffusions building up on Sohl-Dickstein et al. (2015).
@@ 4 Lipschitz constant of the vector field
Under ˜3.3 Lipman et al. (2023), show that the corresponding vector field that generates $p_{t}(\cdot|y)$ for every $y\in\mathbb{R}^{d}$ is given by
@@
A natural first idea is to aim for a global Lipschitz constant, this corresponds to the assumption that $v$ has a bounded Lipschitz constant. The next result shows that such a constant cannot exist in this setting. We can upper and lower bound the Lipschitz constant of $v$:
@@
$B^{t}_{i,j}\coloneqq\Big{\|}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\mathds{1}_{i=j}+\big{(}\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}}\big{)}\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Cov}(Y^{\cdot,t})_{ij}\Big{\|}_{\infty},$
where $Y^{x,t}$ is a random variable with density $q\propto p_{t}(x|\cdot)p^{*}(\cdot)$ for $x\in\mathbb{R}^{d}$. Assume $B_{i,j}$ exists for all $i,j\in\{1,..,d\}$ and define $B\coloneqq(B_{ij})_{i,j=1,...,d}$.
For the Lipschitz constant in space $\Gamma_{t}$ of $v_{t}$ on $\mathbb{R}^{d}$, we have that
@@
The upper and lower bound in Theorem˜4.2 depend on the choice of $\sigma_{t}$. We want to choose $\sigma_{t}$ such that $\big{|}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\big{|}$ does not grow too fast for small $\sigma_{t}$ or large $|\sigma_{t}^{\prime}|$. The next result shows that it is not possible to choose a function $\sigma_{t}$ such that the absolute value of the quotient is smaller than $\log(\sigma_{\min}^{-1})$ for all $t\in[0,1]$ or that the integral over $\big{|}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\big{|}$ remains small.
@@
  1. 1.
    There is a $t^{*}\in[0,1]$ such that
    $\frac{\sigma_{t^{*}}^{\prime}}{\sigma_{t^{*}}}=\log(\sigma_{\min}),$
    for any choice of $\sigma_{t}$ satisfying ˜$3.$
  2. 2.
    For any choice of $\sigma_{t}$ satisfying ˜3.3,
    $\int_{0}^{1}\big{|}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\big{|}\;\mathrm{d}t=\log(\sigma_{\min}^{-1}).$
We note that Lemma˜4.3 does not depend on the specific probability path, it holds for all quotients of smooth choices of $\sigma_{t}$. Furthermore, similar terms arise in all cases, where the chosen probability path implies vector fields of the form (13). The first part of Lemma˜4.3 reveals that a global Lipschitz bound on $v_{t}$ independent of $\sigma_{\min}$ and the covariance term is not feasible in this setting. To compensate for $\frac{\sigma_{t}^{\prime}}{\sigma_{t}}=\log(\sigma_{\min})$ in the case of $i=j$, the second term must have the same absolute value. However, this prohibits a smaller bound when $i\neq j$. The second part of Lemma˜4.3 implies that, even the case of $\operatorname{Cov}(Y^{x,t})_{ij}\lesssim\sigma_{t}^{2}$ for all $i,j$ will lead to a logarithmic dependency on $\sigma_{\min}^{-1}$, which ultimately influences the rate.
To control the Lipschitz constant, we thus need to assume that $p^{*}$ is such that the covariance term in Theorem˜4.2 decays in a controlled order for $\sigma_{t}\rightarrow\sigma_{\min}$ for $i=j$ and the same decay of the off-diagonal elements is fast enough.
@@
$\mathrm{(II)}$ $\operatorname{Var}(Y^{x,t}_{i})$ $=\big{(}\frac{\sigma_{t}}{t^{\gamma}}\big{)}^{2}\big{(}1+O\big{(}\big{(}\frac{\sigma_{t}}{t^{\gamma}}\big{)}^{\frac{1}{\kappa}}\big{)}),$ f $\text{or all }i,$ $t>t^{*},$
@@
Under these assumptions, any choice of $\sigma_{t}$ satisfying ˜3.3 will lead to a bounded Lipschitz constant.
@@
Grant ˜3.3, ˜4.1 and ˜4.4 with fixed parameters $\gamma,\kappa$ and $t^{*}$. Then there is a constant $C$ such that
@@
In order to allow for comparison with other works, we are going to show that the very abstract ˜4.4 is satisfied if the unknown distribution is of the form
@@
This setting was used in Stéphanovitch (2024) to construct Lipschitz continuous pushforward maps in diffusion models.
This setting specifically allows for unbounded distribution that are not $\log$-concave (i.e the logarithmic density is concave) and thus the Brascamp-Lieb inequality (Brascamp and Lieb, 1976, Theorem 4.1) cannot be applied directly. The $\log$-concavity assumption has been popular in the analysis of Diffusion models and Flow Matching, see for example Gao and Zhu (2024); Bruno et al. (2023); Gao et al. (2025, 2024b). First, we note that by Lemma˜7.11 the first assumption in ˜3.2 is satisfied for densities of the form (14) with $\alpha=1$. For simplicity, we choose $\gamma=1$, but the result extends easily to other $\gamma$.
Theorem 4.6.
Set $\gamma=1$. Assume that $\mathbb{P}^{*}$ is of the form (14). Then ˜4.4 is fulfilled with
$C=e^{2L},\quad\kappa=1,\quad\text{ and }\quad t^{*}\quad\text{ such that }\quad\sigma_{t^{*}}=c(L,d),$
where $c(L,d)\in(0,1)$ is a constant that depends only on $L$ and $d$.
The uniform covariance bound is proven using the fact that we can control the effect of bounded perturbations of Gaussians on the variance, which follows from the Holley-Strooke perturbation principle. The bound on the diagonal entries employs proof techniques from the Cramér-Rao lower bound. The bound on the off-diagonal elements is developed from the Brascamp-Lieb type covariance estimate of Menz (2014).
The proof of Theorem˜4.6 reveals that for $t>t^{*}$ large enough and $i\neq j$, the covariance decay is of order $\sigma_{t}^{4}$ and $\kappa=1$. This suggests that the class of distributions for which Theorem˜4.5 applies is significantly larger.

### v2

@@ 2 The Flow Matching objective and Wasserstein evaluation
(Kunkel and Trabs, 2025b, Theorem 3.1) Assume that all functions in $\mathcal{M}$ are Lipschitz continuous for fixed $t$ with Lipschitz constant $\tilde{\Gamma}_{t}.$ Then for any $\tilde{v}\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d},$ with $\tilde{v}\in\mathcal{M}$ and $v$ from (4)
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{\psi}_{1}(Z)})\leq\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi_{1}(Z)})+\sqrt{2e}e^{\int_{0}^{1}\tilde{\Gamma}_{t}\;\mathrm{d}t}\Big(\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ X_{t}\sim p_{t}\end{subarray}}\big[|v_{t}(X_{t})-\tilde{v}_{t}(X_{t})|^{2}\big]\Big)^{1/2}.$ (8)
While this decomposition into the convolution error $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi_{1}(Z)})$ and the error depending on generalization and approximation is natural from a theoretical perspective, it cannot be reversed. As we shall see, different constructions of $v_{t}(\cdot|y)$ and $p_{t}(\cdot|y)$ can lead to the same distribution $\mathbb{P}^{\hat{\psi}_{1}(Z)}$. In case the other construction admits a vector field $\tilde{v}\in\mathcal{M}$, the right hand side of Theorem˜2.1 would be large, while the left hand side is small. We note that the Lipschitz assumption can be relaxed to a one-sided Lipschitz assumption, following the proof of Stéphanovitch (2025, Proposition 1). While this leads to a bound on the largest eigenvalue of $D_{x}v$ instead of the largest entry, we refrain from this generalization to be consistent with previous results and assumptions in the analysis of Flow Matching (Albergo and Vanden-Eijnden, 2023; Benton et al., 2024) and to make the following chapter more intuitive.
The second term in (8) is connected to the suitability of the set $\mathcal{M}$ to approximate $v$. This depends on the specific structure of $v$. The factor $\exp(\int_{0}^{1}2\tilde{\Gamma}_{t}\;\mathrm{d}t)$ results from the use of Grönwall’s inequality and is one of the key difficulties in the analysis of Flow Matching. In the analysis of diffusion models, this can be circumvented using Girsanovs lemma to obtain bounds in the total variation distance. Intuitively, the Brownian motion prevents the exponential propagation of the error. The next example demonstrates this difference when bounding divergences between distributions with underlying dynamics. It also shows that the exponential factor cannot be avoided in a general setting.
Example 2.2.
Let $d=1,\lambda>0$ and $Z\sim\mathcal{N}(0,1)$. Consider a linear vector field $v_{t}(x)=\lambda x$ and a perturbed version $v_{t}^{\varepsilon}(x)=\lambda x+\varepsilon$. Then $\psi_{t}(x)=e^{\lambda t}x$ is the solution of the ODE (3) using $v_{t}$ and $\psi^{\varepsilon}_{t}(x)=e^{\lambda t}x+\frac{\varepsilon}{\lambda}(e^{\lambda t}-1)$ is the solution of the ODE (3) using $v_{t}^{\varepsilon}$. Thus
$\mathsf{W}_{1}(\mathbb{P}^{\psi_{1}(Z)},\mathbb{P}^{\psi^{\varepsilon}_{1}(Z)})=\mathsf{W}_{1}\Big(\mathcal{N}\big(0,e^{2\lambda}\big),\mathcal{N}\Big(\frac{\varepsilon}{\lambda}\big(e^{\lambda t}-1\big),e^{2\lambda}\Big)\Big)=\frac{\varepsilon}{\lambda}\big(e^{\lambda t}-1\big).$
On the other hand, consider two processes $(X^{(1)}_{t})_{t\in[0,T]},(X^{(2)}_{t})_{t\in[0,T]}$ that satisfy
$\mathrm{d}X^{(1)}_{t}=b^{(1)}(X^{(1)}_{t},t)\;\mathrm{d}t+\sigma(t)\;\mathrm{d}B_{t},\quad X^{(1)}_{0}\sim\mathbb{P}$
$\mathrm{d}X^{(2)}_{t}=b^{(2)}(X^{(2)}_{t},t)\;\mathrm{d}t+\sigma(t)\;\mathrm{d}B_{t},\quad X^{(2)}_{0}\sim\mathbb{P}$,
where $(B_{t})_{[0,T]}$ is a Brownian motion. Under suitable assumptions on $\mathbb{P},b^{(1)},b^{(2)}$ and $\sigma$, see Chen et al. (2023c); Oko et al. (2023) and the references therein, one can use Girsanov’s theorem to bound the total variation distance
$\operatorname{TV}(p^{(1)}_{T},p^{(2)}_{T})\lesssim\int_{0}^{T}\mathbb{E}[|b^{(1)}(X^{(1)}_{t},t)-b^{(2)}(X^{(1)}_{t},t)|^{2}]\;\mathrm{d}t.$
In view of Equation˜5, any Lipschitz constraint on the class $\mathcal{M}$ should be aligned with the Lipschitz constant of $v.$ Looking at (4), this Lipschitz constant depends on the construction of $v_{t}(\cdot|\cdot)$ and $p_{t}(\cdot|\cdot)$ and $p^{*}$.
The first term in (8) is determined solely by the conditional vector fields $v_{t}(\cdot|y)$ and is therefore independent of the approximation class $\mathcal{M}$. To bound this term and to analyze the Lipschitz constant in Section˜3, we impose the following assumptions on the densities of the unknown distribution $\mathbb{P}^{*}$ and the latent distribution $\mathbb{U}$.
@@
    where the mean shift $\mu\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ and the variance function $\sigma\colon[0,1]\rightarrow\mathbb{R}_{>0}$ are smooth, component-wise monotone functions such that $\mu_{0}(y)=0,\;\mu_{1}(y)=y,$ and $\sigma_{0}=1,\sigma_{1}=\sigma_{\min},$ for a $\sigma_{\min}>0$.
Under ˜2.3, we know that
see proof of Kunkel and Trabs (2025b, Theorem 3.3), where the first part of the proof does not use the assumption that the support of $\mathbb{P}^{*}$ is compact. Note that the constant in (9) depends only on $\|p^{*}\|_{B^{\alpha}_{1,\infty}}$ and $d$.
˜2.3 means in particular that the latent distribution is $\mathcal{N}(0,I_{d})$. The monotonicity assumption guarantees that there is no unnecessary movement of mass. ˜2.3 coincides with the construction in Lipman et al. (2023, Section 4). Very similar generative models based on probability Flows have been studied by Gao et al. (2024b, a). Additionally, the use of $\mathcal{N}(0,I_{d})$ for the latent distribution is pervasive in generative modeling, e.g. in diffusions building up on Sohl-Dickstein et al. (2015). Often, $\mu$ and $\sigma$ are chosen as linear interpolations between the boundary conditions Lipman et al. (2023); Gao et al. (2024b); Liu et al. (2022), this choice is far from unique, see e.g. the trigonometric interpolation Albergo and Vanden-Eijnden (2023) or the related negative exponential quasi-interpolation in variance preserving diffusion modeling (Song et al., 2021). In case the boundary conditions are met, we recover the same generated distribution $\mathbb{P}^{\psi_{1}(Z)}$, but the corresponding vector fields can differ significantly. In Section˜3, we are going to study the general setting.
@@ 3 Lipschitz constant of the vector field
Motivated by the upper bound in (8), we are going to study the Lipschitz constant of the vector field. Since $v$ is differentiable, our starting point is the Jacobian, which allows for several immediate conclusions. Based on the Jacobian, we derive assumptions on $\mathbb{P}^{*}$ that imply $\int_{0}^{1}\Gamma_{t}\;\mathrm{d}t\leq C$. Afterwards, we study classes of unknown distributions that satisfy these assumptions or that admit no Lipschitz controlled vector field.
Under ˜2.3 Lipman et al. (2023), show that the corresponding vector field that generates $p_{t}(\cdot|y)$ for every $y\in\mathbb{R}^{d}$ is given by
@@
A natural first idea is to aim for a global Lipschitz constant, this corresponds to the assumption that $v$ has a bounded Lipschitz constant. The next result shows that such a constant cannot exist in this setting without considering properties of $\mathbb{P}^{*}$. We can upper and lower bound the Lipschitz constant of $v$.
@@
$B^{t}_{i,j}\coloneqq\sup_{x\in\mathbb{R}^{d}}\Big|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\mathds{1}_{i=j}+\big(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}}\big)\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Cov}(Y^{x,t})_{ij}\Big|_{\infty},$ (11)
where $Y^{x,t}$ is a random variable with density $q\propto p_{t}(x|\cdot)p^{*}(\cdot)$ for $x\in\mathbb{R}^{d}$. Assume $B_{i,j}$ exists for all $i,j\in\{1,..,d\}$ and define $B\coloneqq(B_{ij})_{i,j=1,...,d}$. For the Lipschitz constant in space $\Gamma_{t}$ of $v_{t}$ on $\mathbb{R}^{d}$, we have that
@@
Additionally, there is a $t^{*}\in[0,1]$ such that $\frac{\sigma_{t^{*}}^{\prime}}{\sigma_{t^{*}}}=\log(\sigma_{\min}),$ for any choice of $\sigma_{t}$ satisfying ˜2.3. Hence a global spatial Lipschitz bound on $v_{t}$ independent of $\sigma_{\min}$ and the covariance term is not feasible in this setting.
The distribution with density $q$ is a weighted version of $p^{*}$ which shifts the probability mass to areas closer to $x$ relative to the rest of the mass of $\mathbb{P}^{*}$. The strength of this shift is governed by $\sigma_{t}$, smaller values correspond to a stronger weighing. The covariance depends on both, the covariance of $p^{*}$ and of $\sigma_{t}$. In case $x\in\operatorname{supp}(p^{*})$, $q\overset{\sigma_{\min}\rightarrow 0}{\longrightarrow}\delta_{x}$.
The upper and lower bound in Theorem˜3.2 depend on the choice of $\sigma_{t}$. We want to choose $\sigma_{t}$ such that $\big|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\big|$ does not grow too fast for small $\sigma_{t}$ or large $|\sigma_{t}^{\prime}|$. The next result shows that it is not possible to choose a function $\sigma_{t}$ such that the integral over $\big|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\big|$ remains small.
@@
For any choice of $\sigma_{t}$ satisfying ˜2.3, $\int_{0}^{1}\big|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\big|\;\mathrm{d}t=\log(\sigma_{\min}^{-1}).$
We note that Lemma˜3.3 does not depend on the specific probability path, it holds for all quotients of smooth choices of $\sigma_{t}$. Furthermore, similar terms arise in all cases, where the chosen probability path implies vector fields of the form (10). To compensate for $\frac{\sigma_{t}^{\prime}}{\sigma_{t}}=\log(\sigma_{\min})$ in the case of $i=j$, the second term in (11) must have the same absolute value. However, this prohibits a smaller bound when $i\neq j$. Lemma˜3.3 implies that, even the case of $\operatorname{Cov}(Y^{x,t})_{ij}\lesssim\sigma_{t}^{2}$ for all $i,j$ will lead to a logarithmic dependency on $\sigma_{\min}^{-1}$, which ultimately influences the rate.
To control the Lipschitz constant, we thus need to assume that $p^{*}$ is such that the covariance term in Theorem˜3.2 decays in a controlled order for $\sigma_{t}\rightarrow\sigma_{\min}$ for $i=j$ and the same decay of the off-diagonal elements is fast enough.
@@
$\mathrm{(II)}$ $\operatorname{Var}(Y^{x,t}_{i})$ $=\Big(\frac{\sigma_{t}}{t^{\gamma}}\Big)^{2}\Big(1+O\Big(\Big(\frac{\sigma_{t}}{t^{\gamma}}\Big)^{\frac{1}{\kappa}}\Big)\Big),$ f $\text{or all }i,$ $t>t^{*},$
@@
Under these assumptions, any choice of $\sigma_{t}$ satisfying ˜2.3 will lead to a bounded Lipschitz constant.
@@
Grant ˜2.3, ˜3.1 and ˜3.4 with fixed parameters $\gamma,\kappa$ and $t^{*}$. Then there is a constant $C$ such that
@@
˜3.4 requires careful control over the covariance behavior of the weighted unknown distribution $q$. While the off-diagonal should decay fast enough, the variance must decay at a certain rate. In the following, we present two classes of distributions that satisfy ˜3.4. One of the classes is a subset of the set of $\log$-concave densities, i.e. densities such that the logarithmic density is concave. The other setting shows that ˜3.4 also allows for unbounded distribution that are not $\log$-concave and thus the Brascamp-Lieb inequality (Brascamp and Lieb, 1976, Theorem 4.1) cannot be applied directly. The $\log$-concavity assumption has been popular in the analysis of Diffusion models and Flow Matching, see for example Gao and Zhu (2024); Bruno et al. (2025); Gao et al. (2025, 2024b). Additionally, we show that careful control over the variance is necessary, i.e. if the variance decays too fast, the vector field does not admit a controlled Lipschitz constant.
The first class of distributions are $\log$-concave distributions with $\operatorname{supp}(p^{*})=\mathbb{R}^{d}$ whose potential is three times continuously differentiable and the partial derivatives are bounded, i.e. $p^{*}$ is such that there are $m,L>0$ with
$p^{*}(x)=\exp(-V(x)),\quad V\in C^{3},\quad\|V\|_{C^{3}}\leq L,\quad V\gtrsim mI\quad\forall x\in\mathbb{R}^{d}.$ (12)
In case of a large choice $\sigma_{\min}$, which is not excluded at this point, $L$ should not be too large for technical reasons.
The second class of distributions is of the form
@@
This setting was used in Stéphanovitch (2024) to construct Lipschitz continuous pushforward maps in diffusion models. We note that by Lemma˜A.3 the first assumption in ˜2.3 is satisfied for densities of the form (13) with $\alpha=1$. For simplicity, we choose $\gamma=1$, but the result extends easily to other $\gamma$.
Proposition 3.6.
Set $\gamma=1$. Assume that $\mathbb{P}^{*}$ is of the form (12) or (13). Then ˜3.4 is fulfilled with
[…]
where $c(L,d)\in(0,1)$ is a constant that depends only on $L,d$ and whether $\mathbb{P}^{*}$ is of the form (12) or (13).
In case of Equation˜12, the uniform covariance bound is proved via the Brascamp-Lieb inequality and the behavior of the covariance matrix is shown by a Taylor expansion on $\nabla V$. The proof of the bound for distributions of the form Equation˜13 is mathematically more challenging: The uniform covariance bound is proven using the fact that we can control the effect of bounded perturbations of Gaussians on the variance, which follows from the Holley-Strooke perturbation principle. The bound on the diagonal entries employs proof techniques from the Cramer-Rao lower bound. The bound on the off-diagonal elements is developed from the Brascamp-Lieb type covariance estimate of Menz (2014).
The proof of Proposition˜3.6 reveals that for $t>t^{*}$ large enough and $i\neq j$, the covariance decay is of order $\sigma_{t}^{4}$ and $\kappa=1$. This suggests that the class of distributions for which Theorem˜3.5 applies allows for a slower decaying off-diagonal variance and a less controlled variance.
On the other hand, the next example demonstrates the consequences of a variance that decays too fast, which occurs even in case of a simple $\mathbb{P}^{*}$ such as the one-dimensional uniform distribution. This stresses the importance of a careful Lipschitz analysis.
Example 3.7.
Let $\gamma=1,d=1$ and $\mathbb{P}^{*}=\mathcal{U}([0,1])$. Then $\Gamma_{t}\gtrsim\big|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\big|$.
Example˜3.7 and Lemma˜3.3 show that the exponential term appearing in Theorem˜2.1 cannot be bounded without obtaining a factor $\sigma_{\min}^{-1}$. The proof of Example˜3.7 shows that the derivative of the vector field depends on $\operatorname{Var}_{Y\sim q}(Y)$, which decays arbitrarily fast for large $x$.